Actions
Run Plan 26 June, 2015¶
Tracker alignment with a muon beam, field off.
Daily schedule¶
June | 26 |
estimated time | 10pm - 6am |
MOM | R. Bayes |
Lead shifter | C. Pidcott |
Shifter 2 | A. Dick |
BLOC | C.Pidcott |
SOC | J. Nugent |
Experimenter | |
Physics Shifter | P. Franchini |
Aim | muon beam alignment data |
Beam? | Yes |
Additional Personnel |
Data Requested¶
- Reference Run
- 300 MeV muon beam based alignment data for remaining time. (Currents provided)
Beamline Configurations¶
We expect to be running with 2V beamloss and beam bump active.
Reference Run with Decay Solenoid¶
Please use the run control settings.
Data for Alignment study:¶
Requested: 300 MeV muon beam with 70k TOF2 events. No specific emittance but Ed suggests that the diffuser will reduce the rate too much to be useful. As there are no settings in run control, the following have been taken from the magic spreadsheet.
Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | RunControl tag |
MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | ||
Mu+ | 342 | 338 | 336 | 82 | 85.6 | 106.9 | 74.9 | 261.1 | 558.8 | 131.5 | 244.1 | 300.6 | 199.4 | 204.7 | 309.9 | 264.9 | NA |
- proton absorber setting is unknown but can be taken from the 6-240 MeV+M0 beam in run control.
- ~10000 spills
- If the target number of spills is achieved then go to 6-200+M0 beam and take 70k TOF2 events or 10000 spills ( which ever comes first),
Redo beam line study.¶
- If that is achieved redo the run for Jaroslaw's beam study (which was wiped out by data corruption)
Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | RunControl tag |
MeV/c | MeV/c | MeV/c | mm | A | A | A | A | A | A | A | A | A | A | A | A | ||
Pi+/Mu+ | 408.6 | 405.27 | 238.0 | 82 | 102.4 | 127.9 | 89.0 | 323.2 | 668.6 | 94.2 | 181.07 | 211.53 | 98.54 | 136.50 | 210.49 | 177.76 | 6-200+MatchDown3_With_DS |
- take ~5000 pulses of the above.
Alignment data with diffuse beam.¶
- If that is complete return to 6-200+M0 beam with all irises closed except for iris4 (mask 1110). This should be the diffuser with the thick tungsten iris excluded.
Notes¶
- The 315 A restriction is only for downstream quadrupoles. D1 should be fine at the current given above.
- The solenoid cooling is ongoing. This seemed to work well in parallel with the data collection of the evening of the 25 June. Should this continue it will mean there will be occasional (every 4 hours) controlled access of the hall to change the nitrogen dewars.
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Updated by Bayes, Ryan over 8 years ago ยท 9 revisions